# Memory-Augmented LLM Personalization with Short- and Long-Term Memory Coordination

### v1

@@ Introduction
Retriever $\mathcal{R}$: Retrieval is the process by which the retriever accesses stored knowledge. However, since the memorized knowledge differs between STM and LTM, their retrieval processes also differ. STM is retrieved in the order in which it is stored, while LTM is retrieved through association(e.g., recalling a past mistake by recognizing its similarity(Kahneman 2011)). To address this, we have designed two retrievers: a closest-match retriever, $\mathcal{R}{c}$, for STM retrieval, and a semantic-match retriever, $\mathcal{R}{s}$, for LTM retrieval. $\mathcal{R}_{c}$ aims to find the knowledge stored in STM that is closest to the query in terms of Levenshtein distance, which indicates the minimum number of deletions, insertions, or substitutions required to transform string s into string t (e.g., $lev(^{\prime}test^{\prime},^{\prime}tent^{\prime})=1$ since only one step ${}^{\prime}s^{\prime}->^{\prime}n^{\prime}$ is needed). However, since the retrieval process for LTM is fast and unconscious, we have chosen to train an encoder to obtain semantic embeddings and retrieve knowledge in LTM based on cosine similarity (Madaan et al. 2022).
Data
Data Construction [^2] Data and the construction code will be released here
@@
Existing dialogue datasets often lack awareness of the importance of penalization, while recent works like Xu et al. (2023) have explored the capability of LLMs to generate high-quality chat corpora. In light of this, we propose injecting user profiles into the dialogue generation process using self-chat simulations within real-world conversational scenarios. Specifically, we focus on medical scenarios as they typically involve dialogues between patients and doctors, encompassing a wealth of common-sense information (e.g., Tylenol can alleviate fever) and personal details (e.g., chronic diseases, dialogue preferences). These scenarios allow us to emphasize the significance of memory and penalization, respectively. To obtain personalized dialogues, one straightforward method is to incorporate user’s profile into a language model prompt. We first derive the patient’s profile including personal information, symptoms and dialogue preference from the publicly available medical corpus (Chen et al. 2020) and then follow Xu et al. (2023)’s work using self-chat to guide powerful chat models (e.g., ChatGPT) simulate high-quality dialogues. The difference is that we endow the patient’s and doctor’s profile to the chat model at the beginning of conversation simulation. Further, we prompt the powerful chat model to produce follow-up dialogues related to the same symptom, new symptoms etc. to obtain historical information. The whole construction pipeline and detailed statistics can be seen in Figure 3 and Table 2, respectively.
Safety and Evaluation
@@
Evaluation To further assess our dataset, two master students with medical background are hired for evaluating the quality of the dataset by identifying any dirty content and safety issues on 100 random samples [^3] Quality was scored based on the presence of hallucinations, irrelevant content, etc. Each criterion resulted in a deduction of one point, with a total of 6 points. Safety evaluation focused on identifying profanity, inappropriate suggestions, etc., and any presence of safety issues was indicated by answering ”yes.” For more details, please refer to the data link.. The average quality score was found to be 5.27, and the safety ratio, indicating the proportion of instances without safety issues, was 94%.
Experiments
Setup
To train our model, we leverage the AdamW optimizer(Loshchilov and Hutter 2018) with a learning rate of $5\times 10^{-5}$ and also a linear warm-up scheduler initialized with 10% of the total training steps as warm-up steps and a weight decay of $10^{-4}$ to avoid over-fitting. The LoRA’s rank of update matrices is set as 8 and the scaling factor alpha is 32. To accommodate the task of requirements, we set the maximum length of input and output to 1024 and 2048 tokens, respectively. All implementations are conducted with Pytorch(Paszke et al. 2017), PEFT(Mangrulkar et al. 2022) and Transformers(Wolf et al. 2020) on a computation node configured with a 256G CPU and two 32G Tesla V100 GPUs.
Baselines
Considering the contributions of our work, we opt to compare our DPeM and MaLP with three different configurations: Standard, with dict-based Mem(Madaan et al. 2022) and with LoRA(Hu et al. 2021) in terms of two current SOTA LLMs as base models: GPT3.5 and LLaMA-7B(Touvron et al. 2023).
Tasks and Metrics
We adopt the evaluation methods proposed by Salemi et al. (2023) and Wang et al. (2023) to assess the performance of our personalized LLM model on three tasks:
Question Answering (QA) - We evaluate the effectiveness by posing user-relevant/knowledge-relevant questions to the model and comparing its generated answers with the user profile and memory. The ROUGE-1 and ROUGE-L metrics are used for evaluation;
Preference Classification - We also assess personalization by prompting the trained model to select the user’s dialogue preference from a pre-defined set and measure performance using the Accuracy metric;
Response Generation - In addition to empirical results, we evaluate the quality of responses generated by the trained LLM for new queries from the same user in terms of content and personalization. To do so, we use the scoring method proposed by Wang et al. (2023) and calculate the Win Rate between different settings and standard generation of the base LLM. Furthermore, we conduct human evaluation to validate the alignment of this automatic scoring schema with human judgments.
Comparative Study
ROUGE-1 ROUGE-L ROUGE-1 ROUGE-L Accuracy % Win Rate %
GPT3.5 Standard 32.07 30.81 35.62 31.78 36.31 -
w Mem 34.93 34.27 40.19 38.27 41.73 80.91
w DPeM 40.81 38.78 40.87 39.51 47.72 86.60
LLaMA-7B Standard 21.41 19.82 25.01 23.69 21.42 -
w Mem 21.90 20.44 32.90 31.17 21.15 78.41
w DPeM 22.37 20.97 35.07 33.98 33.06 84.60
w LoRA 30.89 29.66 34.90 33.60 61.05 72.01
w MaLP 35.59 33.91 36.91 36.37 69.95 91.53
Table 3: The main results on different tasks.
Table 3 presents the main evaluation results for Profile/Knowledge QA, Preference Classification, and Response Generation tasks. The addition of memory improves the performance of both GPT3.5 and LLaMA-7b LLMs compared to the standard setting, as it provides additional knowledge prompts to enhance the LLM’s understanding of user queries. However, our novel DPeM exhibits superior performance in assisting LLMs. When combined with GPT3.5 as the base LLM, DPeM outperforms the dict-based memory setting (Madaan et al. 2022) with relative improvements of 13.16% and 3.24% in ROUGE-L scores for profile QA and knowledge QA tasks, respectively. Additionally, DPeM demonstrates better user-specific assistance by achieving a 14.35% increase in classification accuracy compared to dict-based memory and a 7.03% higher win rate for response generation. Similarly, when configured with LLaMA-7b as the base LLM, DPeM achieves relative improvements of 2.59% and 9.02% in profile and knowledge QA tasks, respectively, along with 56.31% and 7.89% enhancements in classifying user preferences and generating personalized responses. These improvements can be attributed to the novel dual-process schema of DPeM, where the rehearsal process refreshes and rewrites knowledge to reduce the risk of retrieving irrelevant information, and the executive process memorizes knowledge in a distinguish-aware manner, leading to more effective retrieval.
One interesting thing we observed is that despite the improvements made by DPeM, it’s still insufficient for acquiring user-specific needs. However, by leveraging PEFT (i.e., LoRA) as can be seen in the results of QA tasks using LLaMA-7b as the base, DPeM achieves a greater improvement on knowledge QA than profile QA while using LoRA achieves a greater improvement on profile QA. Moreover, LoRA helps LLM to know user preference better as it boosts the accuracy of classifying user preference by 39.63% while using DPeM solely improves the accuracy by 11.64% compared with standard setting. However, despite the user-specific need detected by using LoRA, we notice that using LoRA solely is not comparable with using memory on response generation which indicates the importance of memory in our whole MaLP.
By combining DPeM and LoRA into a unified framework, our MaPL approach can effectively incorporate both user-specific needs and knowledge detected from previous dialogue history, resulting in the best performance across all three evaluation tasks compared to other configurations. One more notable thing is that the intricate nature language understanding and generation of LLaMA-7b is not comparable with GPT3.5. However, by using MaLP, LLaMA-7b achieves promising results compared with GPT3.5 towards responding user in a user-favorable manner. These findings further validate the effectiveness and superiority of our novel DPeM mechanism as well as the unified MaLP frame.
Response Quality Study
Case Study
Figure 5: A case study showing the quality of generated response in terms of different settings. Scores are evaluated by human.
Human Judgement
To validate the alignment of our automatic scoring schema with human judgements, we follow the work of Wang et al. (2023) to conduct point-wise evaluation. Specifically, two well-educated master students are hired and 100 response pairs are sampled (i.e., responses generated by standard setting and MaLP using LLaMA-7b). Then we ask the students to indicate which response is better by selecting 1(win), 0(tie) and -1(lose) for each pair. Next, we calculate the Pearson Correlation Coefficient (P.C) and also the accuracy between human scores and automatic scores. The P.C of 0.72 and the accuracy of 84% together indicate the feasibility and high confidence of our evaluation method.
Conclusion & Future Work
In summary, we proposed MaLP which integrates a novel dual-process enhanced memory mechanism and a peft approach to enhance LLMs with awareness of user-specific needs. This straightforward yet effective endeavor enables personalized LLMs while maintaining low resource consumption. Additionally, our innovative data construction method facilitated the creation of high-quality dialogue datasets in the medical domain, and its applicability can be easily extended to other scenarios such as legal consultation, encouraging further exploration in LLM personalization. The extensive experiments and human judgment tests conducted validate the effectiveness of our work.
Despite the success, our method remains in the prototype stage. Next, we will focus on injecting expert knowledge into memory and enabling its online-learning fashion.
Related Work
Memory-Augmented LLM refers to apply a memory that contains user feedback from previous mistakes and by prepending or postpending the new input query with the stored feedback, the output of LLM can be improved(Ouyang et al. 2022). Efforts have been made in terms of the usage of memory. Tandon et al. (2021) first proposed to leverages a corrector that can correct the model’s output in terms of the similar mistake stored in the memory previously. However, this method aims to repair the wrong output while Madaan et al. (2022) argued that the stored experience can be used to avoid incorrect output by prepending/postpending the feedback to the new query. Another usage of memory is to include the memory into a learning frame such as self-learning or teacher-student paradigm so that the LLM can learn by iterative refinement(Madaan et al. 2023; Dalvi, Tafjord, and Clark 2022). In tandem, the key for better usage of memory is to equip powerful retrievers(Guu et al. 2020; Lewis et al. 2020; Yuan et al. 2022) which we argue that memory structure can be improved. The main difference between our work and the previous work is that our work refine the memory structure, instead we design a close-to-real memory mechanism that can better identify and retrieve information for better LLM output and along with PEFT techniques to achieve personalized LLM.
Personalized LLM has seen increasing attentions since it can provide tailored experience that aligns with their user’s expectations in terms of their needs(Salemi et al. 2023). Previous works focused on identifying user preferences by Ceteris Paribus(CP)-nets(Asher, Bonzon, and Lascarides 2010). Unfortunately , this kind of methods suffer from its limited ability of natural language understanding. As LLMs emerged, prompt-based methods attempt to design in-depth prompts such as chain-of-thoughts prompts that can guide LLM to produce desired output with aware of user status and context content(Wang et al. 2023; Wu et al. 2023; Aher, Arriaga, and Kalai 2023). Another way resorts to enhancing LLMs with aware of user information and fine-tuning LLMs to generate responses towards user-specific needs. For example, Korbak et al. (2023); Salemi et al. (2023); Xu et al. (2023) tried to inject user profile information in the pre-training stage and fine-tune the LLM in terms of the learned preferences from user. Unfortunately, fully trained LLMs can be too time- and resource-consuming, thus we propose to leverage parameter-efficient fine-tuning (PEFT) techniques and along with our novel memory mechanism for learning user preferences. In tandem, the most part that differentiates our work with previous work is that we are the first that propose to design a close-to-real memory mechanism instead of a dict-based structure and along with PEFT techniques to achieve and enhance personalized LLMs.
Acknowledgments

### v3

@@ 2.4.3 Components
Retriever $\mathcal{R}$: Retrieval is the process by which the retriever accesses stored knowledge. However, since the memorized knowledge differs between STM and LTM, their retrieval processes also differ. STM is retrieved in the order in which it is stored, while LTM is retrieved through association(e.g., recalling a past mistake by recognizing its similarity(Kahneman, 2011)). To address this, we have designed two retrievers: a closest-match retriever, $\mathcal{R}{c}$, for STM retrieval, and a semantic-match retriever, $\mathcal{R}{s}$, for LTM retrieval. $\mathcal{R}_{c}$ aims to find the knowledge stored in STM that is closest to the query in terms of Levenshtein distance, which indicates the minimum number of deletions, insertions, or substitutions required to transform string s into string t (e.g., $lev(^{\prime}test^{\prime},^{\prime}tent^{\prime})=1$ since only one step ${}^{\prime}s^{\prime}->^{\prime}n^{\prime}$ is needed). However, since the retrieval process for LTM is fast and unconscious, we have chosen to train an encoder to obtain semantic embeddings and retrieve knowledge in LTM based on cosine similarity (Madaan et al., 2022).
3 Data
3.1 Data Construction
Existing dialogue datasets often lack awareness of the importance of penalization, while recent works like Xu et al. (2023) have explored the capability of LLMs to generate high-quality chat corpora. In light of this, we propose injecting user profiles into the dialogue generation process using self-chat simulations within real-world conversational scenarios. Specifically, we focus on medical scenarios as they typically involve dialogues between patients and doctors, encompassing a wealth of common-sense information (e.g., Tylenol can alleviate fever) and personal details (e.g., chronic diseases, dialogue preferences). These scenarios allow us to emphasize the significance of memory and personalization, respectively. To obtain personalized dialogues, one straightforward method is to incorporate user’s profile into a language model prompt. We first derive the patient’s profile including personal information, symptoms and dialogue preference from the publicly available medical corpus [^2] https://github.com/UCSD-AI4H/Medical-Dialogue-System (Chen et al., 2020) and then follow Xu et al. (2023)’s work using self-chat to guide powerful chat models (e.g., ChatGPT) simulate high-quality dialogues. The difference is that we endow the patient’s and doctor’s profile to the chat model at the beginning of conversation simulation. Further, we prompt the powerful chat model to produce follow-up dialogues related to the same symptom, new symptoms etc. to obtain historical information. The whole construction pipeline and detailed statistics can be seen in Figure 3 and Appendix A, respectively.
3.2 Safety and Evaluation
Evaluation To further assess our dataset, two master students with medical background are hired for evaluating the quality of the dataset by identifying any dirty content and safety issues on 100 random samples [^3] Quality scoring is depicted in the Appendix B. The average quality score was found to be 5.27, and the safety ratio, indicating the proportion of instances without safety issues, was 94%.
4 Experiments
4.1 Setup
For the medical knowledge injection, we use the open-source datasets from HealthCareMagic-100k and iCliniq [^4] https://github.com/Kent0n-Li/ChatDoctor (Yunxiang et al., 2023) and set the learning rate as 1e-4, batch size as 20, and weight decay as 0.05 for training. To train our MaLP, we leverage the AdamW optimizer(Loshchilov and Hutter, 2018) with a learning rate of 5e-5 and also a linear warm-up scheduler initialized with 10% of the total training steps as warm-up steps and a weight decay of 1e-4 to avoid over-fitting. The LoRA’s rank of update matrices is set as 8 and the scaling factor alpha is 32. To accommodate the task of requirements, we set the maximum length of input and output to 1024 and 2048 tokens, respectively. All implementations are conducted with Pytorch(Paszke et al., 2017), PEFT(Mangrulkar et al., 2022) and Transformers(Wolf et al., 2020) on a computation node configured with a 256G CPU and two 32G Tesla V100 GPUs.
4.2 Baselines
Considering the contributions of our work, we opt to compare our DPeM and MaLP with three different configurations: Standard, with dict-based Mem(Madaan et al., 2022) and with LoRA(Hu et al., 2021) in terms of three current SOTA LLMs as base models [^5] Due to the resources limitation, we are unable to test larger scale LLMs; Finetuning GPT3.5 is a black-box, we didn’t find a way to apply LoRA on GPT3.5 and some results are omitted. However, the results express the power of MaLP.: GPT3.5, LLaMA-7B, LLaMA-13B(Touvron et al., 2023).
4.3 Tasks and Metrics
We follow the evaluation methods of Salemi et al. (2023) and Wang et al. (2023) to assess the performance of our proposed approach on three tasks:
Question Answering (QA) - We evaluate the effectiveness by posing user-relevant/knowledge-relevant questions to the model and comparing its generated answers with the truth from the user profile and memory. The ROUGE-1 and ROUGE-L metrics are used for evaluation;
Preference Classification - We also assess personalization by prompting the trained model to select the user’s dialogue preference from a pre-defined set and measure performance using Accuracy;
Response Generation - In addition to empirical results, we evaluate the quality of responses generated by the trained LLM for new queries from the same user in terms of the content and preference. To do so, we follow the scoring method of Wang et al. (2023) and calculate the Win Rate between different settings and the standard generation of the base LLM. Furthermore, we conduct human evaluation to validate the alignment of this automatic scoring schema with human judgments.
4.4 Comparative Study
ROUGE-1 ROUGE-L ROUGE-1 ROUGE-L Accuracy % Win Rate %
GPT3.5 Standard 32.07 30.81 35.62 31.78 36.31 -
w Mem 34.93 34.27 40.19 38.27 41.73 80.91
w DPeM 40.81 38.78 40.87 39.51 47.72 86.60
w LoRA - - - - - -
w MaLP - - - - - -
LLaMA-7B Standard 21.41 19.82 25.01 23.69 21.42 -
w Mem 21.90 20.44 32.90 31.17 21.15 78.41
w DPeM 22.37 20.97 35.07 33.98 33.06 84.60
w LoRA 30.89 29.66 34.90 33.60 61.05 72.01
w MaLP 35.59 33.91 36.91 36.37 69.95 91.53
LLaMA-13B Standard 22.67 21.02 26.91 23.98 24.37 -
w Mem 23.10 21.39 34.06 32.47 23.68 78.92
w DPeM 23.57 22.01 36.90 35.09 34.96 84.81
w LoRA 31.29 29.96 36.79 34.99 62.47 71.93
w MaLP 35.97 34.63 37.88 37.07 71.05 91.27
Table 2: The main results on different tasks.
Table 2 presents the main evaluation results for Profile/Knowledge QA, Preference Classification, and Response Generation tasks. The addition of memory improves the performance of both GPT3.5 and LLaMA LLMs compared to the standard setting, as it provides additional knowledge prompts to enhance the LLM’s understanding of user queries. However, our novel DPeM exhibits superior performance in assisting LLMs. When combined with GPT3.5 as the base LLM, DPeM outperforms the dict-based memory setting (Madaan et al., 2022) with relative improvements of 13.16% and 3.24% in ROUGE-L scores for profile QA and knowledge QA tasks, respectively. Additionally, DPeM demonstrates better user-specific assistance by achieving a 14.35% increase in classification accuracy compared to dict-based memory and a 7.03% higher win rate for response generation. Similarly, when configured with LLaMA-7b as the base LLM, DPeM achieves relative improvements of 2.59% and 9.02% in profile and knowledge QA tasks, respectively, along with 56.31% and 7.89% enhancements in classifying user preferences and generating personalized responses. These improvements can be attributed to the novel dual-process schema of DPeM, where the rehearsal process refreshes and rewrites knowledge to reduce the risk of retrieving irrelevant information, and the executive process memorizes knowledge in a distinguish-aware manner, leading to more effective retrieval.
One interesting thing we observed is that despite the improvements made by DPeM, it’s still insufficient for acquiring user-specific needs. However, by leveraging LoRA as can be seen in the results of QA tasks using LLaMA as the base, DPeM achieves a greater improvement on knowledge QA than profile QA while using LoRA achieves a greater improvement on profile QA. Moreover, LoRA helps LLM to know user preference better as it boosts the accuracy of classifying user preference by 39.63% while using DPeM solely improves the accuracy by 11.64% compared with standard setting. However, despite the user-specific need detected by using LoRA, we notice that using LoRA solely is not comparable with using memory on response generation which indicates the importance of memory in our whole MaLP.
By combining DPeM and LoRA into a unified framework, our MaLP approach can effectively incorporate both user-specific needs and knowledge detected from previous dialogue history, resulting in the best performance across all three evaluation tasks compared to other configurations. One more notable thing is that the nuanced distinction in language understanding and generation across various base models may result in subtle differences. These findings further validate the effectiveness and superiority of our novel DPeM mechanism as well as the unified MaLP frame.
4.5 Ablation Study
We further conduct ablation study to validate the completeness of our proposed frame. From table 3, we notice that with knowledge injection, the performance of knowledge QA improves which aligns our intuition to inject domain knowledge first for better responses. When equipped with DPeM, LoRA and fully configured MaLP, the observation stays the same as discussed in the Section 4.4.
ROUGE-1 ROUGE-L ROUGE-1 ROUGE-L Accuracy % Win Rate %
LLaMA-7B Standard 21.41 19.82 25.01 23.69 21.42 -
w Injection 21.39 19.82 33.98 34.11 21.07 73.67
w DPeM 22.37 20.97 35.07 33.98 33.06 84.60
w LoRA 30.89 29.66 34.90 33.60 61.05 72.01
w MaLP 35.59 33.91 36.91 36.37 69.95 91.53
Table 3: The ablation study results on different modules.
4.6 Response Quality Study
4.7 Case Study
Figure 5: A case study showing the quality of generated response in terms of different settings.
4.8 Human Judgement
To validate the alignment of our automatic scoring schema with human judgements, we follow the work of Wang et al. (2023) to conduct point-wise evaluation. Specifically, two master students are hired and 100 response pairs are sampled (i.e., responses generated by standard setting and MaLP using LLaMA-13b). Then we ask the students to indicate which response is better by selecting 1(win), 0(tie) and -1(lose) for each pair. Next, we calculate the Pearson Correlation Coefficient (P.C) and also the accuracy between human scores and automatic scores. The P.C of 0.72 and the accuracy of 84% together indicate the feasibility and high confidence of our evaluation method.
5 Related Work
Memory-Augmented LLM refers to apply a memory that contains user feedback from previous mistakes and by prepending or postpending the new input query with the stored feedback, the output of LLM can be improved(Ouyang et al., 2022). Efforts have been made in terms of the usage of memory. Tandon et al. (2021) first proposed to leverages a corrector that can correct the model’s output in terms of the similar mistake stored in the memory previously. However, this method aims to repair the wrong output while Madaan et al. (2022) argued that the stored experience can be used to avoid incorrect output by prepending/postpending the feedback to the new query. Another usage of memory is to include the memory into a learning frame such as self-learning or teacher-student paradigm so that the LLM can learn by iterative refinement(Madaan et al., 2023; Dalvi et al., 2022). In tandem, the key for better usage of memory is to equip powerful retrievers(Guu et al., 2020; Lewis et al., 2020; Yuan et al., 2022). The main difference between our work and the previous work is that our work refine the memory structure, instead we design a close-to-real memory mechanism that can better identify and retrieve information for enhancement.
Personalized LLM has seen increasing attentions since it can provide tailored experience that aligns with their user’s expectations in terms of their needs(Salemi et al., 2023). Previous works focused on identifying user preferences by Ceteris Paribus(CP)-nets(Asher et al., 2010). Unfortunately , this kind of methods suffer from its limited ability of natural language understanding. As LLMs emerged, prompt-based methods attempt to design in-depth prompts such as chain-of-thoughts prompts that can guide LLM to produce desired output with aware of user status and context content(Wang et al., 2023; Wu et al., 2023; Aher et al., 2023). Another way resorts to enhancing LLMs with aware of user information and fine-tuning LLMs to generate responses towards user-specific needs. For example, Korbak et al. (2023); Salemi et al. (2023); Xu et al. (2023) tried to inject user profile information in the pre-training stage and fine-tune the LLM in terms of the learned preferences from user. Unfortunately, fully trained LLMs can be too resource-consuming, thus we propose to leverage parameter-efficient fine-tuning (PEFT) techniques and along with our novel memory mechanism for personalization. Distinctively, our work stands out from previous research as we pioneer the conception of a realistic memory mechanism and additionally, we employ PEFT techniques to not only attain but also amplify the effectiveness of personalized medical assistant.
6 Conclusion & Future Work
In summary, we proposed MaLP which integrates a novel dual-process enhanced memory mechanism and a peft approach to enhance medical assistants with awareness of user-specific needs. This simple yet effective endeavor enables personalized LLMs while maintaining low resource consumption. Additionally, our innovative data construction method provided the community a fresh perspective to explore personalized medical assistant. The extensive experiments and human judgment tests conducted validate the effectiveness of our work.
Limitations
Despite the empirical success and the production of superior responses, our simple yet effective method remains in the prototype stage. Three notable limitations warrant attention. Firstly, our memory operates in an offline fashion, resembling a smoothly collaborative database. Regrettably, it is incapable of learning from new queries, functioning merely as auxiliary prompts rather than an integral part of the intricate knowledge possessed by the LLM itself. Our dedicated team is actively engaged in the process of incorporating all aspects of memory into the inside of the base LLM. This involves leveraging multiple peft modules to emulate the workflow of the brain’s memory mechanism.
Secondly, the forgetting mechanism in our current implementation relies on frequency counting. However, in scenarios such as avoidance learning (e.g., "fire touch can lead to fire fear"), our DPeM mechanism can encompass a more comprehensive approach. To address this, we plan to introduce learning schemas/losses in the subsequent phase to regulate and control avoidance behavior. We’re excited about making these limitations into novelties in the near future.
Thirdly, applying this technology to real-world scenarios can be complex. For instance, in the case of millions of users, allocating a 7B model for each user could be prohibitively expensive. Alternatively, a large language model (LLM), such as a 175B model, could be employed. This approach involves allocating layers to different users while sharing common community features. However, privacy concerns, such as information leakage, may arise. Our initial approach to addressing this issue is to leverage Federated Learning to model the framework, although further discussion is warranted to delve into the specifics. Our team is dedicated to exploring this direction further.
Ethics Statement
After carefully reviewing the ACL Ethics Policy, we are committed to show our respect and obey to consent all.
Acknowledgements
We gratefully acknowledge support from # NSF-CNS-2154199: Collaborative Research: SaTC: CORE: Medium: Audacity of Exploration: Toward Automated Discovery of Security Flaws in Networked Systems through Intelligent Documentation Analysis.
Appendix A Appendix A. Data Statistics
The statistics of our generated dataset can be seen in the Table 4
@@ Acknowledgements
Appendix B Appendix B. Data Quality Scoring form
@@
Quality was scored based on the presence of hallucinations, irrelevant content, dirty content, invalid symbols, offensive content and harmful suggestions. Each criterion resulted in a deduction of one point, with a total of 6 points. Safety evaluation focused on identifying profanity, inappropriate suggestions and any presence of safety issues was indicated by answering "yes". The scoring table can be seen in Table 5. We calculate the average quality score based on the forms from annotators.
Aspect Explanation Answer
Hallucinations Contains Wrong facts
Invalid symbols Contains invalid symbols
Offensive content Contains insulation / profanity
Dirty content Answers are with unwanted preferences
Harmful suggestions Contains harmful treatment for patients
Irrelevant content The answer is not relevant to the question
Table 5: Quality Scoring Form